# Energetic Electrons Accelerated and Trapped in a Magnetic Bottle above a Solar Flare Arcade

### v1

@@ Abstract
Where and how flares efficiently accelerate charged particles remains an unresolved question. Recent studies revealed that a “magnetic bottle” structure, which forms near the bottom of a large-scale reconnection current sheet above the flare arcade, is an excellent candidate for confining and accelerating charged particles. However, further understanding its role requires linking the various observational signatures to the underlying coupled plasma and particle processes. Here we present the first study combining multi-wavelength observations with data-informed macroscopic magnetohydrodynamics and particle modeling in a realistic eruptive flare geometry. The presence of an above-the-looptop magnetic bottle structure is strongly supported by the observations, which feature not only a local minimum of magnetic field strength but also abruptly slowing down plasma downflows. It also coincides with a compact hard X-ray source and an extended microwave source that bestrides above the flare arcade. Spatially resolved spectral analysis suggests that nonthermal electrons are highly concentrated in this region. Our model returns synthetic emission signatures that are well-matched to the observations. The results suggest that the energetic electrons are strongly trapped in the magnetic bottle region due to turbulence, with only a small fraction managing to escape. The electrons are primarily accelerated by plasma compression and facilitated by a fast-mode termination shock via the Fermi mechanism. Our results provide concrete support for the magnetic bottle as the primary electron acceleration site in eruptive solar flares. They also offer new insights into understanding the previously reported small population of flare-accelerated electrons entering interplanetary space.
@@ 1 Introduction
Figure 1: Formation of a magnetic bottle structure above a solar flare arcade. (a) The presence of a large-scale reconnection current sheet above the flare arcade leads to a local divergence of the magnetic field lines in the above-the-looptop region, forming a magnetic bottle (shaded blue ellipse). The vertical red line is the reconnection current sheet, and the dashed red curve denotes the quasi-separatrix layer. (b) The total magnetic field strength distribution in a 2.5-D resistive MHD model shows a generally lower magnetic field strength in the magnetic bottle region. Also marked are the X point where the field lines break and reconnect, the Y point where the current sheet meets the quasi-separatrix layer, and a fast-mode termination shock formed by reconnection outflows impinging upon the flare arcade.
@@
Previous results based on HXR analysis of the electron time-of-flight distances (Aschwanden et al., 1996a, b) and observations of above-the-looptop (“ALT” hereafter) sources with a high nonthermal electron density (e.g., Krucker et al., 2010; Ishikawa et al., 2011; Krucker & Battaglia, 2014; Fleishman et al., 2022) have suggested that the primary acceleration in large eruptive flares may be located in the cusp region just above the flare arcade, but not necessarily in the upper portion of the current sheet, including the primary X point(s). Recently, by combining microwave imaging spectroscopy observations with magnetohydrodynamics (MHD) modeling, Chen et al. (2020b) found that the region near the bottom of the current sheet (also known as the “Y point” owing to the bifurcation of the current layer; see Figure 1(a)) coincides with a local depression of magnetic field strength. This peculiar structure, referred to as a “magnetic bottle,” is a natural consequence of energy release driven by magnetic reconnection in a large-scale, vertical current sheet above the flare arcade. In the schematic picture shown in Figure 1(a), the anti-parallel magnetic field lines encompassing the current sheet diverge toward the top of the flare arcade. A similar field line divergence is also evident as one follows the footpoints of the flare arcade upward toward the looptop. In turn, the conservation of magnetic flux demands a general reduction of the magnetic field strength in the ALT region. Such a physical picture is supported by analytical models of flare reconnections (Fletcher & Martens, 1998; Lin & Forbes, 2000) and numerical simulations that solve the magnetohydrodynamics (MHD) equations. Figure 1(b) shows an example frame from our MHD simulations (after Shen et al. 2018), which displays a reduced magnetic field strength at the same region.
@@
Figure 2: Microwave and X-ray light curves of the X8.2-class eruptive solar flare on 2017 September 10. The solid vertical line indicates the time of interest of this work during the gradual phase of the flare (around 16:10 UT). The vertical dotted line marks the time studied by Chen et al. (2020b) during the early impulsive phase of the flare when the eruption was initiated, and the double-sided arrow denotes the main impulsive phase of the flare around the primary hard X-ray and microwave peak (see Gary et al. 2018 for more details on the flare evolution).
@@
This magnetic bottle structure coincides with the location of the ALT HXR source and shows a strong concentration of microwave-emitting nonthermal electrons (Chen et al., 2020b), implicating a key role it may be playing in the particle acceleration processes. This is also where the fast reconnection outflows—which carry the bulk of the released energy—collide head-on against the flare arcade to create a plethora of energetic phenomena such as collapsing magnetic traps (Somov & Kosugi, 1997; Karlický & Kosugi, 2004), fast-mode termination shocks (Forbes, 1986; Tsuneta & Naito, 1998; Aurass et al., 2002; Aurass & Mann, 2004; Mann et al., 2009; Guo & Giacalone, 2012; Chen et al., 2015, 2019; Takasao et al., 2015; Polito et al., 2018; Shen et al., 2018, 2022; Ye et al., 2020; Luo et al., 2021; French et al., 2024), slow-mode or gas dynamic shocks (Reeves et al., 2007; Longcope & Guidoni, 2011; Longcope et al., 2016; Longcope & Qiu, 2022), as well as turbulence and oscillations (Takasao & Shibata, 2016; Kontar et al., 2017; Reeves et al., 2020; Ye et al., 2020; Shen et al., 2022, 2023; Ruan et al., 2023; Shibata et al., 2023; Wang et al., 2023b), serving as an ideal environment to heat flare plasma and accelerate charged particles.
@@
However, understanding how the local concentration occurs and, in turn, the underlying particle acceleration and transport mechanisms requires a concerted observational–modeling approach that links the various emission features to the highly coupled plasma dynamics and particle processes, which has been heretofore elusive. Here, we use a novel macroscopic MHD and particle model (after Kong et al. 2019) to produce not only a distribution of thermal plasma in a realistic flare geometry but also a spatially, spectrally, and temporally resolved distribution of energetic electrons throughout the flare region. The model makes it possible, for the first time, to compare the model outputs with multi-wavelength imaging spectroscopy observations that trace both the heated plasma and nonthermal electrons. The paper is organized as follows. Section 2 discusses the multi-wavelength observations and analysis. Section 3 describes the MHD, particle, and emission modeling setup and results. Section 4 put both the observational and modeling results into a coherent physical context and discuss their implications. For the sake of readability, extensive technical details for X-ray and microwave data analysis and numerical modeling are included in the Appendix.
@@
Figure 3: Multi-wavelength observations of the flare during its gradual phase at 16:10 UT. (a) A coronal mass ejection observed in white light by the SOHO/LASCO coronagraph along with a trailing long plasma sheet. Inner inset is the 131 Å EUV channel image made by the Solar Ultraviolet Imager (SUVI) onboard the GOES-R satellite. (b) Detailed view of the plasma sheet and post-flare arcade region. A small box denotes the field of view in (a). The background is the SDO/AIA 131 Å EUV channel image, which samples the hot 10 MK plasma through the Fe XXI line, shown with an inverse color scale (i.e., darker color is brighter). Orange contours are the 7.9 GHz microwave source (15%, 50%, and 90% of the maximum). The RHESSI 25–60 keV hard X-ray image is shown as red contours (10% and 50% of the maximum). (c) Same as (b), but the background is the Fe XXIV line EUV image observed by Hinode/EIS, which is sensitive to 18 MK plasma. (d) Similar to (b), but with filled contours denoting multi-frequency EOVSA microwave images from 3.4 GHz to 12.9 GHz.
@@ 2 Observations
The observations were obtained during the well-observed X8.2-class eruptive solar flare on 2017 September 10, recorded by ground- and space-based instruments at multiple wavelengths. This event, thanks to its favorable viewing perspective, has a geometry that matches very well the standard model of eruptive solar flares (Figure 3). We refer interested readers to Chen et al. (2020a) and references therein for a more detailed discussion on its three-dimensional (3D) configuration based on multi-wavelength, multi-perspective observations. Briefly, the event is induced by an erupting magnetic flux rope that drives a fast white light coronal mass ejection (CME). Immediately trailing the CME core, a long linear-shaped feature, seen in both white light and EUV, extends down to the top of the bright flare arcade, interpreted as a large-scale reconnection current sheet viewed edge-on (Warren et al., 2018; Yan et al., 2018; Chen et al., 2020b). The flare arcade anchored at the solar surface displays a cusp shape as observed by the Atmospheric Imaging Assembly onboard the Solar Dynamics Observatory (SDO/AIA; Lemen et al. 2012) and the EUV Imaging Spectrometer (EIS; Culhane et al. 2007) onboard Hinode (see Figures 3(b) and (c), respectively). The cusp-shaped flare loops are the signature of highly bent magnetic field lines resulting from ongoing magnetic reconnection in the current sheet, producing fast sunward reconnection outflows (e.g., Longcope et al., 2018; Hayes et al., 2019; Yu et al., 2020).
In order to carry out the comparison with our combined MHD, particle, and emission model (see Section 3), we select a period around 16:10 UT during the gradual phase of the eruptive solar flare when the eruption has already propagated to a large distance (). As shown in Figure 3(a), the core of the white light CME is located at $\sim$4 solar radii, followed by a long trailing plasma sheet. In comparison, the size of the flare arcade itself is only $\sim$20 Mm, or $\sim$1of the length of the sheet. Therefore, the plasma processes near the solar surface can be well approximated by a system driven by an infinitely long current sheet as depicted in Figure 1, which allows us to perform detailed data–model comparison in a realistic flare geometry.
We note that the flare event occurred on the west solar limb, and the eruption generally proceeded along the east-west direction. For the sake of easier data-model comparison, throughout this paper, we have rotated the observed images by 90 degrees counter-clockwise and assigned $x=950^{\prime\prime}$ and $y=-140^{\prime\prime}$ in the original Helioprojective Cartesian coordinates as the origin of our new coordinate system. The length units are in megameters, with $1^{\prime\prime}\approx 0.73$ Mm at a distance of 1 AU.
@@ 2.1 X-ray and Microwave Imaging and Spectroscopy
In the X-ray images obtained by the Reuven Ramaty High Energy Solar Spectroscopic Imager (RHESSI; Lin et al. 2002), this period features an ALT HXR source at 25–60 keV (Figure 3(b); see also Figure 11 in Appendix for images at additional energy bands). At the top of the EUV flare arcade, a brighter X-ray source is present (referred to as the “looptop” source), which is dominated by thermal bremsstrahlung from hot ($>$15 MK) plasma with a high density ($\sim\!10^{12}\,\mathrm{cm}^{-3}$). In the total-flux X-ray spectrum, at above $\sim$30 keV, a nonthermal component dominates the spectrum with a power-law shape (Figure 4). Spectral analysis suggests that if this component falls into the thin-target bremsstrahlung regime, it corresponds to a total nonthermal electron density of $n_{e}^{>50}\approx 4\times 10^{6}$ cm-3 above 50 keV (see detailed discussions in Appendix Section A).
@@
Figure 4: RHESSI X-ray spectral analysis results for the time of interest at 16:10 UT. (a) Black curve represents the observed X-ray photon spectrum averaged over 16:10:00 UT and 16:11:08 UT. The red curve is the best-fit spectrum that includes two isothermal components and one nonthermal component arising from thin-target bremsstrahlung from a power-law electron distribution. The two thermal components, shown as the green and yellow curves, have temperatures of 15 MK and 27 MK, and volume emission measures of $1.6\times 10^{51}\,\mathrm{cm}^{-3}$ and $4.1\times 10^{49}\,\mathrm{cm}^{-3}$, respectively. The nonthermal thin-target component has a normalization factor of $1.6\times 10^{54}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$ with a power-law index of $\delta^{\rm thin}=4.7$.
@@
Figure 5: Observed and modeled multi-frequency microwave images from 3.4 GHz to 10.4 GHz. Left panels: Microwave images observed by EOVSA on 2017 September 10 at 16:10:36 UT. Right panels: Synthetic EOVSA microwave images generated from the combined MHD and particle model. Overlaid gray curves are magnetic field lines derived from the MHD model.
@@
Figure 6: Spatially resolved microwave spectra and derived source parameters along the above-the-looptop, arcade-like microwave source. (a) Frequency–space spectrogram obtained along a fiducial slit drawn along the spine of the microwave source (white dashed curve in Figure 3(d)). (b) Example microwave brightness temperature spectra at five sampled locations along the slit (vertical dashed lines in (a) marked from 1 to 5). Black circles with error bars are the measurements, and blue curves are a subset of 200 randomly selected model spectra from the MCMC runs, each of which has a total of 800,000 samples. (c) and (d) Distribution of best-fit magnetic field strength $B$ and total nonthermal electron density above 50 keV $n_{e}^{>50}$ along the slit. The double-side blue arrows mark the approximate location of the magnetic bottle region.
@@
In the microwave images observed by the Expanded Owens Valley Solar Array (EOVSA; Gary et al. 2018) in 2.5–18 GHz, no counterpart of the bright looptop X-ray source is present. Instead, it features an arcade-shaped source that bestrides above the bright EUV flare arcade (orange contours in Figure 3). The arcade-shaped source can be clearly seen in microwave images at all frequencies, with subtle changes in their appearance (left panels in Figure 5). Interestingly, the low-frequency images ($\lesssim$6 GHz) display a small “gap” at the central location (i.e., near $x=0$ Mm) with a relatively lower brightness temperature. With EOVSA’s multi-frequency imaging capability, one can derive microwave spectra from different locations of the images and perform spectral analysis. In Figure 6(a), we show a frequency–space spectrogram obtained along a fiducial slit drawn along the spine of the microwave arcade passing the ALT HXR source at $x\approx 0$ Mm (dashed curve in Figure 3(d)). Such a spectrogram is akin to those obtained by slit spectrographs, with the microwave intensity along the slit “dispersed” in frequency over the vertical axis. Figure 6(b) show microwave brightness temperature spectra derived from five equally spaced locations along the slit. All spectra show a positive spectral slope at low frequencies and a negative slope at high frequencies with a peak brightness temperature of $\sim$100 MK, consistent with gyrosynchrotron emission produced by nonthermal electrons gyrating in the coronal magnetic field (e.g., Dulk & Marsh, 1982). The spectral properties are primarily sensitive to parameters of the nonthermal electron distribution and the magnetic field in the source.
@@
Following Chen et al. (2020b), we use the Markov chain Monte Carlo (MCMC) method to analyze the observed spectra (see Appendix Section B for details). Figure 6(c) shows the distribution of the best-fit magnetic field strength $B$ along the slit. It displays a local minimum of $B$ at the center of the slit around $x\approx 0$ Mm, suggestive of a magnetic bottle structure co-located with the ALT HXR source. Meanwhile, despite the relatively large uncertainties, the distribution of the total nonthermal electron density above 50 keV $n_{e}^{>50}$ (Figure 6(d)) indicates that the best-fit value in the central magnetic bottle region is two orders of magnitude greater than locations into the loop legs at $\lvert x\rvert>20$ Mm, implying that microwave-emitting nonthermal electrons are strongly concentrated in the magnetic bottle region. The best-fit nonthermal electron density in the magnetic bottle region is also consistent with that returned from the X-ray spectral analysis within uncertainties.
@@
Figure 7: Plasma outflows slow down and transform to a cusp shape as they pass the magnetic bottle hosting the Y point. (a) Context SDO/AIA 171 Å EUV image at 16:06:09 UT. (b)–(e) Time-series of running-ratio SDO/AIA 171 Å images. Each plot shown is a ratio of the current image to a time 36-s earlier. The arrows indicate the downward plasma outflow that transforms from a linear shape to a cusp shape as it passes the Y point. (f) Corresponding MHD model showing the vertical component of the flow velocity. Downward plasma outflows (blue color) dominate the region above the Y point. (g) Time-distance plot of SDO/AIA 171 Å running-ratio time-series images obtained along a vertical slice at the central current sheet region (white dashed line in (a)). Approximate tracks of the bottom of the erupting flux rope, Y point, and looptop are marked as dashed curves. The example outflow event in (b)-(e) is marked by the two white dashed lines, which display an abrupt speed change after it passes the Y point. (h) Same as (g), but showing the original SDO/AIA 171 Å intensity.

### v2

@@ Abstract
Where and how flares efficiently accelerate charged particles remains an unresolved question. Recent studies revealed that a “magnetic bottle” structure, which forms near the bottom of a large-scale reconnection current sheet above the flare arcade, is an excellent candidate for confining and accelerating charged particles. However, further understanding its role requires linking the various observational signatures to the underlying coupled plasma and particle processes. Here we present the first study combining multiwavelength observations with data-informed macroscopic magnetohydrodynamics and particle modeling in a realistic eruptive flare geometry. The presence of an above-the-loop-top magnetic bottle structure is strongly supported by the observations, which feature not only a local minimum of magnetic field strength but also abruptly slowing down plasma downflows. It also coincides with a compact hard X-ray source and an extended microwave source that bestrides above the flare arcade. Spatially resolved spectral analysis suggests that nonthermal electrons are highly concentrated in this region. Our model returns synthetic emission signatures that are well matched to the observations. The results suggest that the energetic electrons are strongly trapped in the magnetic bottle region due to turbulence, with only a small fraction managing to escape. The electrons are primarily accelerated by plasma compression and facilitated by a fast-mode termination shock via the Fermi mechanism. Our results provide concrete support for the magnetic bottle as the primary electron acceleration site in eruptive solar flares. They also offer new insights into understanding the previously reported small population of flare-accelerated electrons entering interplanetary space.
@@ 1 Introduction
Figure 1: Formation of a magnetic bottle structure above a solar flare arcade. (a) The presence of a large-scale reconnection current sheet above the flare arcade leads to a local divergence of the magnetic field lines in the above-the-loop-top region, forming a magnetic bottle (shaded blue ellipse). The vertical red line is the reconnection current sheet, and the dashed red curve denotes the quasi-separatrix layer. (b) The total magnetic field strength distribution in a 2.5D resistive MHD model shows a generally lower magnetic field strength in the magnetic bottle region. Also marked are the X point where the field lines break and reconnect, the Y point where the current sheet meets the quasi-separatrix layer, and a fast-mode termination shock formed by reconnection outflows impinging upon the flare arcade.
@@
Previous results based on HXR analysis of the electron time-of-flight distances (Aschwanden et al., 1996a, b) and observations of above-the-loop-top (hereafter ALT) sources with a high nonthermal electron density (e.g., Krucker et al., 2010; Ishikawa et al., 2011; Krucker & Battaglia, 2014; Fleishman et al., 2022) have suggested that the primary acceleration in large eruptive flares may be located in the cusp region just above the flare arcade but not necessarily in the upper portion of the current sheet, including the primary X point(s). Recently, by combining microwave imaging spectroscopy observations with magnetohydrodynamics (MHD) modeling, Chen et al. (2020b) found that the region near the bottom of the current sheet (also known as the “Y point” owing to the bifurcation of the current layer; see Figure 1(a)) coincides with a local depression of magnetic field strength. This peculiar structure, referred to as a “magnetic bottle,” is a natural consequence of energy release driven by magnetic reconnection in a large-scale, vertical current sheet above the flare arcade. In the schematic picture shown in Figure 1(a), the antiparallel magnetic field lines encompassing the current sheet diverge toward the top of the flare arcade. A similar field line divergence is also evident as one follows the footpoints of the flare arcade upward toward the loop top. In turn, the conservation of magnetic flux demands a general reduction of the magnetic field strength in the ALT region. Such a physical picture is supported by analytical models of flare reconnections (Fletcher & Martens, 1998; Lin & Forbes, 2000) and numerical simulations that solve the MHD equations. Figure 1(b) shows an example frame from our MHD simulations (after Shen et al. 2018), which displays a reduced magnetic field strength at the same region.
@@
Figure 2: Microwave and X-ray light curves of the X8.2-class eruptive solar flare on 2017 September 10. The solid vertical line indicates the time of interest of this work during the gradual phase of the flare (around 16:10 UT). The dotted vertical line marks the time studied by Chen et al. (2020b) during the early impulsive phase of the flare when the eruption was initiated, and the double-sided arrow denotes the main impulsive phase of the flare around the primary HXR and microwave peak (see Gary et al. 2018 for more details on the flare evolution).
@@
This magnetic bottle structure coincides with the location of the ALT HXR source and shows a strong concentration of microwave-emitting nonthermal electrons (Chen et al., 2020b), implicating a key role it may be playing in the particle acceleration processes. This is also where the fast reconnection outflows—which carry the bulk of the released energy—collide head-on against the flare arcade to create a plethora of energetic phenomena such as collapsing magnetic traps (Somov & Kosugi, 1997; Karlický & Kosugi, 2004), fast-mode termination shocks (Forbes, 1986; Tsuneta & Naito, 1998; Aurass et al., 2002; Aurass & Mann, 2004; Mann et al., 2009; Guo & Giacalone, 2012; Chen et al., 2015, 2019; Takasao et al., 2015; Polito et al., 2018; Shen et al., 2018, 2022; Ye et al., 2020; Luo et al., 2021; French et al., 2024), slow-mode or gas dynamic shocks (Reeves et al., 2007; Longcope & Guidoni, 2011; Longcope et al., 2016; Longcope & Qiu, 2022), and turbulence and oscillations (Takasao & Shibata, 2016; Kontar et al., 2017; Reeves et al., 2020; Ye et al., 2020; Shen et al., 2022, 2023; Ruan et al., 2023; Shibata et al., 2023; Wang et al., 2023b), serving as an ideal environment to heat flare plasma and accelerate charged particles.
@@
However, understanding how the local concentration occurs and, in turn, the underlying particle acceleration and transport mechanisms requires a concerted observational-modeling approach that links the various emission features to the highly coupled plasma dynamics and particle processes, which has been heretofore elusive. Here, we use a novel macroscopic MHD and particle model (after Kong et al. 2019) to produce not only a distribution of thermal plasma in a realistic flare geometry but also a spatially, spectrally, and temporally resolved distribution of energetic electrons throughout the flare region. The model makes it possible, for the first time, to compare the model outputs with multiwavelength imaging spectroscopy observations that trace both the heated plasma and nonthermal electrons. The paper is organized as follows. Section 2 discusses the multiwavelength observations and analysis. Section 3 describes the MHD, particle, and emission modeling setup and results. Section 4 puts both the observational and modeling results into a coherent physical context and discusses their implications. For the sake of readability, extensive technical details for X-ray and microwave data analysis and numerical modeling are included in the Appendices.
@@
Figure 3: Multiwavelength observations of the flare during its gradual phase at 16:10 UT. (a) A CME observed in white light by the SOHO/LASCO coronagraph along with a trailing long plasma sheet. The inset is the 131 Å EUV channel image made by the Solar Ultraviolet Imager (SUVI) on board the GOES-R satellite. (b) Detailed view of the plasma sheet and postflare arcade region. A small box denotes the field of view in (a). The background is the SDO/AIA 131 Å EUV channel image, which samples the hot 10 MK plasma through the Fe XXI line, shown with an inverse color scale (i.e., darker color is brighter). Orange contours are the 7.9 GHz microwave source (15%, 50%, and 90% of the maximum). The RHESSI 25–60 keV HXR image is shown as red contours (10% and 50% of the maximum). (c) Same as (b), but the background is the Fe XXIV line EUV image observed by Hinode/EIS, which is sensitive to 18 MK plasma. (d) Similar to (b) but with filled contours denoting multifrequency EOVSA microwave images from 3.4 to 12.9 GHz.
@@ 2 Observations
The observations were obtained during the well-observed X8.2-class eruptive solar flare on 2017 September 10, recorded by ground- and space-based instruments at multiple wavelengths. This event, thanks to its favorable viewing perspective, has a geometry that matches very well the standard model of eruptive solar flares (Figure 3). We refer interested readers to Chen et al. (2020a) and references therein for a more detailed discussion on its three-dimensional (3D) configuration based on multiwavelength, multiperspective observations. Briefly, the event is induced by an erupting magnetic flux rope that drives a fast white-light coronal mass ejection (CME). Immediately trailing the CME core, a long, linear feature, seen in both white light and extreme ultraviolet (EUV), extends down to the top of the bright flare arcade, interpreted as a large-scale reconnection current sheet viewed edge-on (Warren et al., 2018; Yan et al., 2018; Chen et al., 2020b). The flare arcade anchored at the solar surface displays a cusp shape as observed by the Atmospheric Imaging Assembly (AIA; Lemen et al. 2012) on board the Solar Dynamics Observatory (SDO) and the EUV Imaging Spectrometer (EIS; Culhane et al. 2007) on board Hinode (see Figures 3(b) and (c), respectively). The cusp-shaped flare loops are the signature of highly bent magnetic field lines resulting from ongoing magnetic reconnection in the current sheet, producing fast sunward reconnection outflows (e.g., Longcope et al., 2018; Hayes et al., 2019; Yu et al., 2020).
In order to carry out the comparison with our combined MHD, particle, and emission model (see Section 3), we select a period around 16:10 UT during the gradual phase of the eruptive solar flare when the eruption has already propagated to a large distance (i.e., 12 minutes after the main microwave and HXR peak at around 15:58 UT; Figure 2). As shown in Figure 3(a), the core of the white-light CME is located at $\sim$4 $R_{\odot}$, followed by a long trailing plasma sheet. In comparison, the size of the flare arcade itself is only $\sim$20 Mm, or $\sim$1of the length of the sheet. Therefore, the plasma processes near the solar surface can be well approximated by a system driven by an infinitely long current sheet as depicted in Figure 1, which allows us to perform detailed data–model comparison in a realistic flare geometry.
We note that the flare event occurred on the west solar limb, and the eruption generally proceeded along the east–west direction. For the sake of easier data–model comparison, throughout this paper, we have rotated the observed images by 90 counter-clockwise and assigned $x=950^{\prime\prime}$ and $y=-140^{\prime\prime}$ in the original helioprojective Cartesian coordinates as the origin of our new coordinate system. The length units are in megameters, with $1^{\prime\prime}\approx 0.73$ Mm at a distance of 1 AU.
@@ 2.1 X-ray and Microwave Imaging and Spectroscopy
In the X-ray images obtained by the Reuven Ramaty High Energy Solar Spectroscopic Imager (RHESSI; Lin et al. 2002), this period features an ALT HXR source at 25–60 keV (Figure 3(b); see also Figure 11 in Appendix for images at additional energy bands). At the top of the EUV flare arcade, a brighter X-ray source is present (referred to as the “loop-top” source), which is dominated by thermal bremsstrahlung from hot ($>$15 MK) plasma with a high density ($\sim\!10^{12}\,\mathrm{cm}^{-3}$). In the total-flux X-ray spectrum, at above $\sim$30 keV, a nonthermal component dominates the spectrum with a power-law shape (Figure 4). Spectral analysis suggests that if this component falls into the thin-target bremsstrahlung regime, it corresponds to a total nonthermal electron density of $n_{e}^{>50}\approx 4\times 10^{6}$ cm-3 above 50 keV (see detailed discussions in Appendix A).
@@
Figure 4: RHESSI X-ray spectral analysis results for the time of interest at 16:10 UT. (a) The black curve represents the observed X-ray photon spectrum averaged over 16:10:00 UT and 16:11:08 UT. The red curve is the best-fit spectrum that includes two isothermal components and one nonthermal component arising from thin-target bremsstrahlung from a power-law electron distribution. The two thermal components, shown as the green and yellow curves, have temperatures of 15 MK and 27 MK and volume emission measures of $1.6\times 10^{51}\,\mathrm{cm}^{-3}$ and $4.1\times 10^{49}\,\mathrm{cm}^{-3}$, respectively. The nonthermal thin-target component has a normalization factor of $1.6\times 10^{54}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$ with a power-law index of $\delta^{\rm thin}=4.7$.
@@
Figure 5: Observed and modeled multifrequency microwave images from 3.4 to 10.4 GHz. Left panels: microwave images observed by EOVSA on 2017 September 10 at 16:10:36 UT. Right panels: synthetic EOVSA microwave images generated from the combined MHD and particle model. Overlaid gray curves are magnetic field lines derived from the MHD model.
@@
Figure 6: Spatially resolved microwave spectra and derived source parameters along the above-the-loop-top, arcade-like microwave source. (a) Frequency–space spectrogram obtained along a fiducial slit drawn along the spine of the microwave source (white dashed curve in Figure 3(d)). (b) Example microwave brightness temperature spectra at five sampled locations along the slit (vertical dashed lines in (a) marked from 1 to 5). Black circles with error bars are the measurements, and blue curves are a subset of 200 randomly selected model spectra from the MCMC runs, each of which has a total of 800,000 samples. (c) and (d) Distribution of the best-fit magnetic field strength $B$ and total nonthermal electron density above 50 keV $n_{e}^{>50}$ along the slit. The double-sided blue arrows mark the approximate location of the magnetic bottle region.
@@
In the microwave images observed by the Expanded Owens Valley Solar Array (EOVSA; Gary et al. 2018) in 2.5–18 GHz, no counterpart of the bright loop-top X-ray source is present. Instead, it features an arcade-shaped source that bestrides above the bright EUV flare arcade (orange contours in Figure 3). The arcade-shaped source can be clearly seen in microwave images at all frequencies, with subtle changes in their appearance (left panels in Figure 5). Interestingly, the low-frequency images ($\lesssim$6 GHz) display a small “gap” at the central location (i.e., near $x=0$ Mm) with a relatively lower brightness temperature. With EOVSA’s multifrequency imaging capability, one can derive microwave spectra from different locations of the images and perform spectral analysis. In Figure 6(a), we show a frequency–space spectrogram obtained along a fiducial slit drawn along the spine of the microwave arcade passing the ALT HXR source at $x\approx 0$ Mm (dashed curve in Figure 3(d)). Such a spectrogram is akin to those obtained by slit spectrographs, with the microwave intensity along the slit “dispersed” in frequency over the vertical axis. Figure 6(b) shows microwave brightness temperature spectra derived from five equally spaced locations along the slit. All spectra show a positive spectral slope at low frequencies and a negative slope at high frequencies with a peak brightness temperature of $\sim$100 MK, consistent with gyrosynchrotron emission produced by nonthermal electrons gyrating in the coronal magnetic field (e.g., Dulk & Marsh, 1982). The spectral properties are primarily sensitive to the parameters of the nonthermal electron distribution and the magnetic field in the source.
@@
Following Chen et al. (2020b), we use the Markov Chain Monte Carlo (MCMC) method to analyze the observed spectra (see Appendix B for details). Figure 6(c) shows the distribution of the best-fit magnetic field strength $B$ along the slit. It displays a local minimum of $B$ at the center of the slit around $x\approx 0$ Mm, suggestive of a magnetic bottle structure colocated with the ALT HXR source. Meanwhile, despite the relatively large uncertainties, the distribution of the total nonthermal electron density above 50 keV $n_{e}^{>50}$ (Figure 6(d)) indicates that the best-fit value in the central magnetic bottle region is 2 orders of magnitude greater than locations in the loop legs at $\lvert x\rvert>20$ Mm, implying that microwave-emitting nonthermal electrons are strongly concentrated in the magnetic bottle region. The best-fit nonthermal electron density in the magnetic bottle region is also consistent with that returned from the X-ray spectral analysis within uncertainties.
@@
Figure 7: Plasma outflows slow down and transform to a cusp shape as they pass the magnetic bottle hosting the Y point. (a) Context SDO/AIA 171 Å EUV image at 16:06:09 UT. (b)–(e) Time series of running-ratio SDO/AIA 171 Å images. Each plot shown is a ratio of the current image to a time 36 s earlier. The arrows indicate the downward plasma outflow that transforms from a linear shape to a cusp shape as it passes the Y point. (f) Corresponding MHD model showing the vertical component of the flow velocity. Downward plasma outflows (blue color) dominate the region above the Y point. (g) Time–distance plot of SDO/AIA 171 Å running-ratio time-series images obtained along a vertical slice at the central current sheet region (white dashed line in (a)). Approximate tracks of the bottom of the erupting flux rope, Y point, and loop top are marked as dashed curves. The example outflow event in (b)-(e) is marked by the two white dashed lines, which display an abrupt speed change after it passes the Y point. (h) Same as (g) but showing the original SDO/AIA 171 Å intensity.